OpenStudy (anonymous):

solve: 3^4x-7 = 4^2x+3

sam (.sam.):

3^(4x-7)=4^(2x+3) ln(3^(4x-7))=ln(4^(2x+3)) 4xln(3)-7ln(3)=ln(4^(2x+3)) 4xln(3)-7ln(3)-2xln(4)=3ln(4) 2x(2ln(3)-ln(4))=7ln(3)+3ln(4) \[2x=\frac{7\ln(3)+3\ln(4)}{2\ln(3)-\ln(4)}\] \[x=\frac{7\ln(3)+3\ln(4)}{2(2\ln(3)-\ln(4))}\]

sam (.sam.):

x=0.6166

OpenStudy (anonymous):

thank you. so you used ln to solve it, but i dont understand how you substituted or set up the equation.

myininaya (myininaya):

@evisabello nothing was substituted ... He took ln( ) of both sides and used the property \[\ln(y^r)=r \ln(y) \] Then he used the distributive property a(b+c)=ab+ac on both sides! Then he put all of his terms with the factor x on one side and all the terms without the factor x on the other side. Then the side with terms that included x, he factored x out and divided by the junk next to the x on both sides.